OpenStudy (anonymous):

Use the vector product to find the area of the triangle PQR, P = (5,1,1), Q = (2,−1,3) and R = (−1,4,2).

OpenStudy (anonymous):

do you know what the answer is already? i got 1.5, not sure if what im going is right though. Doesnt sound right lol

OpenStudy (anonymous):

find vector pq and pr take the dot product of the two the cross product will give you the volume

OpenStudy (anonymous):

its 12.1 or so.... but im stuck on the easiest bit...

OpenStudy (anonymous):

well not the easiest, just getting the vectors pq and pr

OpenStudy (anonymous):

find the magnitude of the cross product between PQ and PR. This will give you the area of the parallelogram. Divide this by 2 and you will get the area of the triangle.

OpenStudy (anonymous):

you would subtract the coordinates of p from q to get vector pq, same for pr

OpenStudy (anonymous):

=1/2((pq.pq)(pr.pr)-(pq.pr)^2)^1/2 =1/2(17x46-14^2^1/2)=12.1

OpenStudy (anonymous):

oh ok well the solutions i have are wrong then for this question;.

OpenStudy (anonymous):

partly wrong... .thanks guys/